(brand) 37. Pituluny dependen CHAPTER 9 Drugs Used in Hormonal, Endocrine, and Reproductive Disorders 193 1 A 1200-lb Holstei I would like to compare my answers. im not sure if I did them right: questions 41.- 46.
Answer #1

41.

1ml = 50 mcg

X = 100 mcg

Therefore X = 100/50 = 2

So dose to be administered is 2 ml

​​​​​​42.

1 pound = 0.4535 kg

1100 pound = 498.9512 kg

Dose to be given daily for 1100lb mare = 0.044*498.9512 = 21.95mg/day

Each ml of Altrenogest 0.22% contains 2.2mg of Altrenogest

Dose to be administered= 21.95mg/2.2 mg= 9.97= 10 ml

43.

1 lb= 0.4535 kg

50lb = 0.4535* 50 = 22.6796kg

Dose to be administered= 0.02mg/kg

For 22.6796kg= 0.02*22.6796= 0.4535

=0.5 mg when rounded off

Soloxine available dose is 0.5 mg/tablet

Therefore 1 tablet should be given with each dose.

44.

1 lb = 0.4535 kg

To convert 13 lb to kg= 13*0.4535= 5.896 kg

= 6 when rounded off

Dose to be administered/kg = 0.5 units

For 6 kg = 0.5*6 = 3

3 units should be administered with each dose

45.

Dose to be administered/lb = 3 mg

Dose for 20 lb poodle= 20*3 = 60mg/day

So, 1 tablet to be administered daily

Hence 14 tablets should be dispensed for 2 weeks

46.

Felimazole 2.5 mg (i.e) 1 tablet will be given for each dose

So 2 tablets are needed for a day.

Dose to be dispensed for 3 weeks = 2* 21 days = 42 tablets

42 tablets should be dispensed for 3 weeks